OpenStudy (anonymous):

explain how the exponent of 10 changes when you multiply a number written in scientific notation by 100. show an example.

OpenStudy (australopithecus):

so just say we had 1*10^2 = 1*(10)(10) = 100 if we increase the exponent by 1 we get 1*10^3 = 1*(10)(10)(10) = 1000 if we decrease the exponent by 5 we get \[1*10^{-2}= \frac{1}{10*10} = \frac{1}{100} = 0.01\] if we decrease by 1 more we get \[1*10^{-3}= \frac{1}{10*10*10} = \frac{1}{1000} = 0.001\] yeah I hope this helps

OpenStudy (australopithecus):

so if we had (1.0*10^2)*100 = 1.0*(10)(10)(100) = 10000.0 = 1.0 * 10^4